# Cameras as Relative Positional Encoding

### v1

@@ Abstract
Transformers are increasingly prevalent for multiview computer vision tasks, where geometric relationships between viewpoints are critical for 3D perception. To leverage these relationships, multiview transformers must use camera geometry to ground visual tokens in 3D space. In this work, we compare techniques for conditioning transformers on cameras: token-level raymap encodings, attention-level relative pose encodings, and a new relative encoding we propose—Projective Positional Encoding (PRoPE)—that captures complete camera frustums, both intrinsics and extrinsics, as a relative positional encoding. Our experiments begin by showing how relative camera conditioning improves performance in feedforward novel view synthesis, with further gains from PRoPE. This holds across settings: scenes with both shared and varying intrinsics, when combining token- and attention-level conditioning, and for generalization to inputs with out-of-distribution sequence lengths and camera intrinsics. We then verify that these benefits persist for different tasks, stereo depth estimation and discriminative spatial cognition, as well as larger model sizes. Code is available on our project webpage [^1] https://www.liruilong.cn/prope/.
@@ 1 Introduction
Images of our world exist in the context of the viewpoints they were captured from. The geometry of these viewpoints—intrinsic and extrinsic parameters that give pixel coordinates their physical meaning—ground visual observations in 3D space. This spatial grounding is increasingly important, especially as advances in 3D vision and embodied intelligence make multiview tasks more ubiquitous.
@@
To solve multiview tasks with transformers, models must bind viewpoint information to patch tokens from each input image. This binding requires special care: just as naive positional encoding techniques for 1D sequences can hinder performance for learning in language models [1], naive encodings of camera geometry may also be suboptimal for multiview vision models [2, 3, 4]. Advances in both settings can be summarized as transitions from absolute [5] to relative encodings [6].
@@
In this work, we study the problem of conditioning vision transformers on the camera geometry of input images. We survey existing techniques for addressing this, which include (i) absolute encodings in the form of pixel-aligned, token-level raymaps—these are the most common in recent state-of-the-art models [7, 8, 9, 10]—and (ii) attention-level relative encodings based on SE(3) pose relationships [3, 4]. We then present a new camera conditioning technique, Projective Positional Encoding (PRoPE), that is designed to capture the complete geometry of cameras as a relative positional encoding. PRoPE models viewing frustum relationships that describe both intrinsics and extrinsics, while remaining easy to incorporate with standard transformer architectures and fused attention kernels [11].
@@
Our main experiments are on three tasks, which span six datasets. We begin with a series of studies comparing camera conditioning techniques for feedforward novel view synthesis using RealEstate10K [12] and Objaverse [13]. Our results highlight the advantages of relative encodings—particularly PRoPE—compared to absolute ones. We then verify that these benefits extend to other settings: we demonstrate improvements when integrating PRoPE into UniMatch [14] for stereo depth estimation across three benchmarks, for a discriminative spatial cognition task using DL3DV [15], and when scaling to larger novel view synthesis models [7, 8].
Our contributions are as follows:
  1. 1.
@@
    Survey. We survey both absolute raymap and relative SE(3) conditioning techniques for camera geometry in multiview transformers.
  2. 2.
@@
    Method. We propose PRoPE (Projective Positional Encoding), a new relative positional encoding technique that injects complete camera geometry—both intrinsics and extrinsics—into a transformer’s self-attention blocks.
  3. 3.
@@
    Evaluation. We present a series of novel view synthesis (NVS) experiments that compare camera conditioning techniques empirically. Our results confirm the advantages of relative pose encoding methods like CAPE [4] and GTA [3], while demonstrating further improvements from PRoPE across a range of settings: scenes with shared intrinsics, scenes with varying intrinsics, for hybrid conditioning that combines both token-level and attention-level representations, and in generalization for out-of-distribution test inputs.
  4. 4.
@@
    Task generalization. We show that the benefits of cameras as relative positional encoding generalize (i) to stereo depth estimation when integrated into UniMatch, (ii) to discriminative spatial cognition, and (iii) when scaling to larger model sizes.
@@
Figure 1: Cameras as relative positional encoding. Language models and multiview transformers must both bind “positional” information to input tokens, in terms of sequence position for language and camera parameters for multiview computer vision. We present a study on camera conditioning that includes absolute positional encodings (raymaps), relative pose encodings [2, 4], and a new method captures relative projective relationships between more complete camera frustums.
@@ 2 Related Work
Absolute and relative positional encodings.
Transformer architectures are permutation-invariant; they therefore require explicit position encoding to understand token order in sequential inputs [5]. Approaches for position encoding in sequence models has been an active area of research [1, 16, 17, 18, 19]. While early works [20, 21, 22, 23, 24, 25, 26, 27] focused on absolute positional encoding (APE), recent methods have increasingly adopted relative positional encoding (RPE), particularly RoPE [28], as a standard across domains, including natural language processing [29, 30, 1, 31] and computer vision [32, 33, 34, 28, 35]. Relative encodings aim to improve models by defining positions as relative offsets between token pairs. These offsets are injected into the pairwise interactions of standard dot product attention [5]:
@@
where $Q,K,V\in\mathbb{R}^{T\times d}$. Position offsets can be injected using the pairwise nature of the $QK^{\top}\in\mathbb{R}^{T\times T}$ matrix, either via additive biases [6, 36, 37] or SO(2)-based rotation [1]. RPE offers important advantages over APE, including translation invariance, improved relationship modeling, and generalization to long sequences [38, 1, 39]. In this work, we study both absolute and relative encodings for conditioning transformers on camera geometry instead of 1D position.
Multiview transformers.
@@
Many computer vision tasks are multiview—they take multiple images and known camera geometry for each image as input. Examples exist in 3D reconstruction and view synthesis [40, 41, 42, 43, 44], pose estimation [45], depth prediction [14, 46, 47], 3D scene understanding [48, 49], robotics [50], and world models [51]. Many recent works leverage the improved scaling properties [22, 52, 53] of vision transformers for solving these tasks [8, 7, 9, 10, 54]. These models slice input images into patches, and use each patch as an independent visual token for the transformer. In this work, we use the model designs proposed by LVSM [8] and UniMatch [14] as a starting point for studying a critical design decision—how transformers are conditioned on camera geometry.
Camera conditioning in transformers.
@@
The dominant approach for conditioning multiview transformers on cameras is currently raymaps [45, 7, 8, 10]—per-pixel 6D embeddings that contain either ray origins and directions [44, 7] or Plücker coordinates [55, 45]. Concatenating these parameters to pixels enables conditioning on both camera intrinsics and extrinsics at the token level. Raymaps, however, require defining a frame of reference [56, 7, 57], which is problematic because the choice of world coordinate system is arbitrary and can hinder generalization. While this problem can be partially addressed by normalizing camera poses [58, 57, 8], prior works have shown that a more fundamental fix is possible through relative encoding [2, 3, 4]. Notably, attention-level encodings that capture relative SE(3) poses don’t require defining a consistent global frame, are compatible with fused attention kerenels [11], and have been shown to improve novel view synthesis performance [3, 4]. In the following section, we survey both absolute raymap and relative SE(3) methods for encoding camera geometry for transformers. We then propose a new relative encoding method, PRoPE, that represents relationships between more complete camera frustums as a relative positional encoding.
3 Conditioning Transformers on Camera
In this work, we evaluate techniques for conditioning multiview vision transformers on camera geometry. To describe these methods, we begin by reviewing the projection matrices used to define camera viewpoints. We then describe their connection to existing geometry encoding techniques: both absolute, token-level encodings in the form of raymaps and relative, attention-level encodings of SE(3) pose. We then present PRoPE: a camera conditioning technique built on GTA [3] that
@@ 3.1 Preliminaries
In this work, we study transformers that take $N$ images from known cameras as input:
@@
where each $\boldsymbol{I}_{i}\in\mathbb{R}^{H\times W\times 3}$ is an image, $\boldsymbol{K}_{i}\in\mathbb{R}^{3\times 3}$ is the camera intrinsics, and $\boldsymbol{T}^{cw}_{i}=(\mathbf{R}^{\text{cw}}_{i},\mathbf{t}^{\text{cw}}_{i})\in\text{SE}(3)$ is the transformation for computing camera coordinates from world coordinates. The latter two terms encode the viewing frustum that corresponds to each image: intrinsics capture the shape and field-of-view of the frustum, while extrinsics capture position and orientation. These are encapsulated in the “world-to-image” projection matrix $\boldsymbol{P}_{i}\in\mathbb{R}^{3\times 4}$:
@@
For notational convenience, these $3\times 4$ projection matrices can be made invertible by lifting to $4\times 4$ with the standard basis vector $\mathbf{e}_{4}=(0,0,0,1)^{\top}$:
$\boldsymbol{\tilde{P}}_{i}=\begin{bmatrix}\boldsymbol{P}_{i}\\ \mathbf{e}_{4}^{\top}\end{bmatrix}.$ (4)
This transformation maps 3D world coordinates to a projective image space defined by the frustum of camera $i$. It can be used to compute 2D image coordinates from world coordinates:
$\begin{bmatrix}\tilde{\mathbf{x}}_{i}\\ 1\end{bmatrix}\propto\tilde{\boldsymbol{P}}_{i}\tilde{\mathbf{X}}_{\text{world}},$ (5)
@@
3.2 Pixel-aligned Geometry Encoding
Token-level, pixel-aligned raymaps are the dominant method for encoding geometry in multiview transformers [7, 42, 10, 9]. Networks that employ raymaps concatenate images $\boldsymbol{I}_{i}\in\mathbb{R}^{H\times W\times 3}$ with per-pixel raymaps $\mathbf{M}_{i}\in\mathbb{R}^{H\times W\times R}$ along the channel dimension, which expands inputs to $\mathbb{R}^{H\times W\times(3+R)}$. There are two main approaches for computing these raymaps, which we refer to as “naive” and Plücker.
Naive raymaps. Naive raymaps [7] are computed as per-pixel origin and direction vectors:
@@ 3.2 Pixel-aligned Geometry Encoding
Plücker raymaps. Plücker raymaps [45, 8] can be implemented by replacing the origin term in naive raymaps with a moment term:
@@
Properties. Raymaps offer a simple approach for conditioning on both camera intrinsics and extrinsics. An important drawback, however, is that they are absolute: similar to early position encoding techniques for 1D sequences [5], raymaps are expressed in global terms. They are therefore sensitive to the arbitrary choice of reference frame, which can hinder generalization.
@@ 3.3 Relative SE(3) Encoding
To remove the need for a global reference frame, recent works have introduced relative encodings for SE(3) camera poses. Two existing approaches fall under this category: CaPE [4] and GTA [3]. Given images $i_{1}$ and $i_{2}$, both aim to condition networks on $\boldsymbol{T}^{cw}_{i_{1}}(\boldsymbol{T}^{cw}_{i_{2}})^{-1}$ using modified self-attention blocks. This captures dense relationships—how each camera pose is situated relative to every other camera pose—and makes networks invariant to how the world frame $w$ is defined.
@@
CaPE [4]. CaPE injects relative SE(3) pose by transforming the $Q$ and $K$ matrices before they are passed to self-attention. CaPE can be formalized using per-token block-diagonal matrices $\mathbf{D}^{\texttt{CaPE}}_{t}$, which is computed by diagonally repeating the camera extrinsics:
@@
Like RoPE [1], transformations are then applied to the $Q$ and $K$ matrices before self-attention. This can be encapsulated into an augmented self-attention block:
@@
GTA [3]. GTA proposes a formulation for per-token transformations with a similar high-level goal as CaPE. GTA’s attention variant transforms the $Q$ and $K$ matrices the same way, while proposing to also transform the $V$ matrix:
@@
where $\alpha_{t_{1},t_{2}}$ is a softmax score computed from the transformed dot product (Equation 14). GTA’s experiments [3] compare several SE(3)-based formulations for $\mathbf{D}^{\texttt{GTA}}$, with and without the value matrix transformation. The best-performing methods include the value transform, and both SE(3) for camera pose and RoPE [1] for 2D patch position. Our experiments include GTA using these terms.

### v2

@@ Abstract
Transformers are increasingly prevalent for multiview computer vision tasks, where geometric relationships between viewpoints are critical for 3D perception. To leverage these relationships, multiview transformers must use camera geometry to ground visual tokens in 3D space. In this work, we compare techniques for conditioning transformers on cameras: token-level raymap encodings, attention-level relative pose encodings, and a new relative encoding—Projective Positional Encoding (PRoPE)—that captures complete camera frustums, both intrinsics and extrinsics, as a relative positional encoding. Our experiments begin by showing how relative conditioning methods improve performance in feedforward novel view synthesis, with further gains from PRoPE. This holds across settings: scenes with both shared and varying intrinsics, when combining token- and attention-level conditioning, and for generalization to inputs with out-of-distribution sequence lengths and camera intrinsics. We then verify that these benefits persist for different tasks, stereo depth estimation and discriminative spatial cognition, as well as larger model sizes. Code is available on our project webpage [^1] https://www.liruilong.cn/prope/.
@@ 1 Introduction
Images of our world exist in the context of the viewpoints they were captured from. The geometry of these viewpoints—intrinsic and extrinsic parameters that give pixel coordinates their physical meaning—ground visual observations in 3D space. This spatial grounding is increasingly important in deep learning, especially as advances in 3D vision and embodied intelligence make multiview tasks more ubiquitous.
@@
To solve multiview tasks with transformers, models must bind viewpoint information to patch tokens from each input image. This binding requires special care: just as naive positional encoding techniques for 1D sequences hinder performance for learning in language models [su2024roformer], naive encodings of camera geometry may also be suboptimal for multiview vision models [safin2023repast, miyato2023gta, kong2024eschernet]. Advances in both settings can be summarized as transitions from absolute [vaswani2017attention] to relative [shaw2018self] encodings.
@@
In this work, we study the problem of conditioning vision transformers on the camera geometry of input images. We survey existing techniques for addressing this, which include (i) absolute encodings in the form of pixel-aligned, token-level raymaps—these are the most common in recent state-of-the-art models [gao2024cat3d, jin2024lvsm, zhou2025stable, weber2025fillerbuster]—and (ii) attention-level relative encodings based on SE(3) pose relationships [miyato2023gta, kong2024eschernet]. We then present a new camera conditioning technique, Projective Positional Encoding (PRoPE), that is designed to capture the complete geometry of cameras as a relative positional encoding. PRoPE models viewing frustum relationships that describe both intrinsics and extrinsics, while remaining easy to incorporate with standard transformer architectures and fused attention kernels [dao2023flashattention].
@@
Our experiments are on three tasks, which span six datasets. We begin with a series of studies comparing camera conditioning techniques for feedforward novel view synthesis using RealEstate10K [zhou2018stereo] and Objaverse [deitke2023objaverse]. Our results highlight the advantages of relative encodings—particularly PRoPE—compared to absolute ones. We then verify that these benefits extend to other settings: we demonstrate improvements when integrating PRoPE into UniMatch [xu2023unifying] for stereo depth estimation across three benchmarks, for a discriminative spatial cognition task using DL3DV [ling2024dl3dv], and when scaling to larger novel view synthesis models [gao2024cat3d, jin2024lvsm].
The contributions of this paper are as follows:
@@
(1) Survey. We survey both absolute raymap and relative SE(3) conditioning techniques for camera geometry in multiview transformers.
@@
(2) Method. We propose PRoPE (Projective Positional Encoding), a new relative positional encoding technique that injects both camera intrinsics and extrinsics into a transformer’s self-attention blocks.
@@
(3) Evaluation. We present a series of novel view synthesis (NVS) experiments that compare camera conditioning techniques empirically. Our results highlight the advantages of relative pose encoding methods like CAPE [kong2024eschernet] and GTA [miyato2023gta], while demonstrating further improvements from PRoPE across a range of settings: scenes with shared intrinsics, scenes with varying intrinsics, for hybrid conditioning that combines both token-level and attention-level representations, and in generalization for out-of-distribution test inputs.
@@
(4) Task generalization. We show that the benefits of cameras as relative positional encoding generalize (i) to stereo depth estimation when integrated into UniMatch, (ii) to discriminative spatial cognition, and (iii) when scaling to larger model sizes.
@@
Figure 1: Cameras as relative positional encoding. Language models and multiview transformers must both bind “positional” information to input tokens, in terms of sequence position for language and camera parameters for multiview computer vision. We present a study on camera conditioning that includes absolute positional encodings (raymaps), relative pose encodings [safin2023repast, kong2024eschernet], and a new method captures relative projective relationships between more complete camera frustums.
@@ 2 Related Work
Absolute and relative positional encodings. Transformer architectures are permutation-invariant; they therefore require explicit position encoding to understand token order in sequential inputs [vaswani2017attention]. Position encoding in sequence models has been an active area of research [su2024roformer, kazemnejad2023impact, peng2023yarn, golovneva2024contextual, schenck2025learning]. While early works [radford2018improving, devlin2019bert, dosovitskiy2021image, caron2021emerging, radford2021learning, rombach2022high, zhai2023sigmoid, kirillov2023segment] focused on absolute positional encoding (APE), recent methods have increasingly adopted relative positional encoding (RPE), particularly RoPE [heo2024rotary], as a standard across domains, including natural language processing [raffel2020exploring, touvron2023llama, su2024roformer, guo2025deepseek] and computer vision [liu2023visual, esser2024scaling, flux2024, heo2024rotary, genmo2024mochi]. Relative encodings aim to improve models by defining positions as relative offsets between token pairs. These offsets are injected into the pairwise interactions of standard dot product attention [vaswani2017attention]:
@@
where $Q,K,V\in\mathbb{R}^{T\times d}$. Position offsets can be injected using the pairwise nature of the $QK^{\top}\in\mathbb{R}^{T\times T}$ matrix, either via additive biases [shaw2018self, raffel2020t5, press2022alibi] or SO(2)-based rotation [su2024roformer]. RPE offers important advantages over APE, including translation invariance, improved relationship modeling, and generalization to long sequences [press2021train, su2024roformer, dubey2024llama]. In this work, we study both absolute and relative encodings for conditioning transformers on camera geometry instead of 1D position.
@@
Multiview transformers. Many computer vision tasks are multiview: they take multiple images and known camera geometry for each image as input. Examples exist in 3D reconstruction and view synthesis [suhail2022light, sajjadi2022scene, wu2024cat4d, liang2024feed, mildenhall2021nerf], pose estimation [zhang2024cameras], depth prediction [xu2023unifying, yang2024depth, ke2024repurposing], 3D scene understanding [hong20233d, zhu2024llava], robotics [shridhar2023perceiver], and world models [po2025long]. Many recent works leverage the improved scaling properties [dosovitskiy2021image, liu2021swin, peebles2023scalable] of vision transformers for solving these tasks [jin2024lvsm, gao2024cat3d, zhou2025stable, weber2025fillerbuster, bar2025navigation]. These models slice input images into patches, and use each patch as an independent visual token for the transformer. In this work, we use the model designs proposed by LVSM [jin2024lvsm] and UniMatch [xu2023unifying] as a starting point for studying a critical design decision—how transformers are conditioned on camera geometry.
@@
Camera conditioning in transformers. The dominant approach for conditioning multiview transformers on cameras is currently raymaps [zhang2024cameras, gao2024cat3d, jin2024lvsm, weber2025fillerbuster]: per-pixel 6D embeddings that contain either ray origins and directions [mildenhall2021nerf, gao2024cat3d] or Plücker coordinates [plucker1865xvii, zhang2024cameras]. Concatenating these parameters to pixels enables conditioning on both camera intrinsics and extrinsics at the token level. Raymaps, however, require defining a frame of reference [mildenhall2019local, gao2024cat3d, guizilini2025zero], which is problematic because the choice of world coordinate system is arbitrary and can hinder generalization. While this problem can be partially addressed by normalizing poses [kulhanek2022viewformer, guizilini2025zero, jin2024lvsm], prior works have shown that a more fundamental fix is possible through relative parameters [safin2023repast, miyato2023gta, kong2024eschernet, yi2025egoallo]. Notably, attention-level encodings that capture relative SE(3) poses don’t require defining a consistent global frame, are compatible with fused attention kernels [dao2023flashattention], and have been shown to improve novel view synthesis performance [miyato2023gta, kong2024eschernet]. In the following section, we survey both absolute raymap and relative SE(3) methods for encoding camera geometry for transformers. We then propose a new relative encoding method, PRoPE, that encodes relationships between more complete camera frustums.
3 Conditioning Transformers on Cameras
@@ 3.1 Preliminaries
We study transformers that take $N$ images from known cameras as input:
@@
where each $\boldsymbol{I}_{i}\in\mathbb{R}^{H\times W\times 3}$ is an image, $\boldsymbol{K}_{i}\in\mathbb{R}^{3\times 3}$ is the camera intrinsics, and $\boldsymbol{T}^{cw}_{i}=(\mathbf{R}^{\text{cw}}_{i},\mathbf{t}^{\text{cw}}_{i})\in\text{SE}(3)$ is the transformation for computing camera coordinates from world coordinates. The latter two terms encode the viewing frustum that corresponds to each image: intrinsics capture the shape and field-of-view of the frustum, while extrinsics capture position and orientation. Both intrinsics and extrinsics are encapsulated in the “world-to-image” projection matrix $\boldsymbol{P}_{i}\in\mathbb{R}^{3\times 4}$:
@@
For notational convenience, these $3\times 4$ projection matrices can be made invertible by lifting to $4\times 4$ with the standard basis vector $\mathbf{e}_{4}=(0,0,0,1)^{\top}$. This transformation maps 3D world coordinates to a projective image space defined by the frustum of camera $i$. It can be used to compute 2D image coordinates from world coordinates:
$\boldsymbol{\tilde{P}}_{i}=\begin{bmatrix}\boldsymbol{P}_{i}\\ \mathbf{e}_{4}^{\top}\end{bmatrix};\quad\begin{bmatrix}\tilde{\mathbf{x}}_{i}\\ 1\end{bmatrix}\propto\tilde{\boldsymbol{P}}_{i}\tilde{\mathbf{X}}_{\text{world}},$ (4)
@@
3.2 Pixel-aligned Camera Encoding
Token-level, pixel-aligned raymaps are the dominant method for encoding geometry in multiview transformers [gao2024cat3d, wu2024cat4d, weber2025fillerbuster, zhou2025stable]. Networks that employ raymaps concatenate images $\boldsymbol{I}_{i}\in\mathbb{R}^{H\times W\times 3}$ with per-pixel raymaps $\mathbf{M}_{i}\in\mathbb{R}^{H\times W\times R}$ along the channel dimension, which expands inputs to $\mathbb{R}^{H\times W\times(3+R)}$. There are two main approaches for computing these raymaps, which we refer to as “naive” and Plücker.
Naive raymaps. Naive raymaps [gao2024cat3d] are computed as per-pixel origin and direction vectors:
@@ 3.2 Pixel-aligned Camera Encoding
Plücker raymaps. Plücker raymaps [zhang2024cameras, jin2024lvsm] can be implemented by replacing the origin term in naive raymaps with a moment term:
@@
Properties. Raymaps offer a simple approach for conditioning on both camera intrinsics and extrinsics. An important drawback, however, is that they are absolute: similar to early position encoding techniques for 1D sequences [vaswani2017attention], raymaps are expressed in global terms. They are sensitive to the arbitrary choice of reference frame, which can hinder generalization.
@@ 3.3 Relative SE(3) Encoding
To remove the need for a global reference frame, recent works have introduced relative encodings for SE(3) camera poses. Two existing approaches fall under this category: CaPE [kong2024eschernet] and GTA [miyato2023gta]. Given images $i_{1}$ and $i_{2}$, both aim to condition networks on $\boldsymbol{T}^{cw}_{i_{1}}(\boldsymbol{T}^{cw}_{i_{2}})^{-1}$ using modified self-attention blocks. This captures dense relationships—how each camera pose is situated relative to every other camera pose—and makes networks invariant to how the world frame $w$ is defined.
@@
CaPE [kong2024eschernet]. CaPE injects relative SE(3) pose by transforming the $Q$ and $K$ matrices before they are passed to self-attention. CaPE can be formalized using per-token block-diagonal matrices $\mathbf{D}^{\texttt{CaPE}}_{t}$, which is computed by diagonally repeating the camera extrinsics:
@@
Like RoPE [su2024roformer], transformations are then applied to the $Q$ and $K$ matrices before self-attention. This can be encapsulated into an augmented self-attention block:
@@
GTA [miyato2023gta]. GTA proposes a formulation for per-token transformations with a similar high-level goal as CaPE. GTA’s attention variant transforms the $Q$ and $K$ matrices the same way, while proposing to also transform the $V$ matrix:
@@
where $\alpha_{t_{1},t_{2}}$ is a softmax score computed from the transformed dot product (Equation 13). GTA’s experiments [miyato2023gta] compare several SE(3)-based formulations for $\mathbf{D}^{\texttt{GTA}}$, with and without the value matrix transformation. The best-performing methods include the value transform, and both SE(3) for camera pose and RoPE [su2024roformer] for 2D patch position. Our experiments include GTA using these terms.